How do I know if the quadratic equation has two real solutions, a double root, or no real solution, without solving the equation in a discriminant way? here are some examples
2a2+3a+1=0
2)2a(a - 3) - 3=3
3a2 - 2a(a - 2)2 - 4=0

Answer by josgarithmetic(39617)

Why without discriminant?
Could you factorize your equations?


Your last equation is not quadratic.
3a^2-2a(a-2)^2-4=0
3a^2-2a(a^2-4a+4)-4=0
3a^2-2a^3+8a^2-8a-4=0
-2a^3+11a^2-8a-4=0
3a^3-11a^2+8a+4=0